OpenStudy (sasogeek):

I just went out to buy a 1.5l bottle of coke. if I drink 300ml every 11minutes, how long will it take me to finish drinking everything in the bottle?

Parth (parthkohli):

Is this a trick question?

OpenStudy (sasogeek):

nope, i really did go out to buy the drink and i want to know how long it will last if I drink at that rate. :P

Parth (parthkohli):

\[1500:x=300:11 \]There's a nice proportion. lol

OpenStudy (lgbasallote):

long enough to get you to the toilets

OpenStudy (sasogeek):

hahaha, well idk if i'm really going to finish it, seeing as there's a fuss everywhere that coke isn't good lol i may end up not drinking everything :P

OpenStudy (anonymous):

\[t =\frac{ 11min }{300 ml }\ x \ 1.5L \frac{ 1000ml }{1L} = 16500 min = 275hrs\]

OpenStudy (sasogeek):

WHAAAAAAAAAAAAT!?

OpenStudy (anonymous):

almost two weeks

OpenStudy (lgbasallote):

lol @cguy \[t = \frac{1500}{300} \times 11\]